In 1943, during world war II, copper was needed for the war effort, so pennies were made of lead. A lead penny had a diameter of 1.8 cm and a thickness of 0.1 cm. If lead has a mass of 11.4 grams per cm^3, what was the mass of a lead penny?

Found 2 solutions by ankor@dixie-net.com, mathman409:
Answer by ankor@dixie-net.com(12701)

In 1943, during world war II, copper was needed for the war effort, so pennies were made of lead. A lead penny had a diameter of 1.8 cm and a thickness of 0.1 cm. If lead has a mass of 11.4 grams per cm^3, what was the mass of a lead penny?
:
A penny is in the shape of a cylinder so we can use the that formula:
V = pi*r^2*h, in our problem: r = 1.8/2 = .9 cm; h = .1 cm
:
V = pi * .9^2 * .1
:
V = pi * .81 * .1
:
V = .254469 cu cm
:
Mass = vol * density:
:
Mass = .254469 * 11.4
:
Mass = 2.9 grams

Answer by mathman409(67)

:
A penny is in the shape of a cylinder so we can use the that formula:
V = pi*r^2*h, in our problem: r = 1.8/2 = .9 cm; h = .1 cm
:
V = pi * .9^2 * .1
:
V = pi * .81 * .1
:
V = .254469 cu cm
:
Mass = vol * density:
:
Mass = .254469 * 11.4
:
Mass = 2.9 grams